What is the molarity of a solution containing 15.0g NaOH?

(molar mass =40.0g?mol) per 250.0mL of solution?
Answer #1

Molarity = no. of moles of solute / volume in L

no. of moles of solute = 15 / 40 = 0.375 moles

volume = 250 mL = 0.25 L

Molarity = 0.375 / 0.25 = 1.5 M